# Finiteness of totally geodesic hypersurfaces

### v2

@@ 1. Introduction
In the above statement, “hyperbolic” refers to a Riemannian manifold with constant strictly negative sectional curvature. Equivalently, up to rescaling the Riemannian metric, $M$ is isometric to $\mathbb{H}^{n}/\Gamma$ where $\mathbb{H}^{n}$ is hyperbolic $n$-space, and $\Gamma$ is a lattice which is moreover arithmetic (see [Mar91, Ch. IX] for the notion of arithmetic lattice).
@@
    Our proof works under the more general condition that the geodesic flow of $(M,g)$ is Anosov, see e.g. [Ebe73, Thm. 3.2] for equivalent characterizations of this property. Our proof also applies to $M$ of finite volume instead of compact, with some additional technical assumptions spelled out in § 2.1.2 and § 2.1.7.
@@
    The arithmeticity of $M$ is a consequence of the prior conclusions and [BFMS21, Thm. 1.1]. See also [MM22, Thm. 1.1] for the case $n=3$.
@@
    Arithmetic hyperbolic manifolds of the first type have infinitely many closed totally geodesic immersed hypersurfaces. All other arithmetic hyperbolic manifolds do not. See [BBKS21, Corollary 5.12] and the remark following [BBKS21, Theorem 1.2].
@@
See [Car51] and [Daj90, Thm. 1.8] for modern presentations of the proof. The axiom of planes, which requires that through any three points in a Riemannian manifold passes a totally geodesic surface, was originally introduced by Riemann [Rie54].
Theorem 2 is not only a significant motivation for our work but also a key step in our proof. We note that in this context there is no version of Theorem 1 for manifolds of positive curvature or for general non-compact manifolds, see § 4.1. We also remark that there is always an open dense set of metrics on a manifold for which there are no totally geodesic submanifolds of dimension greater than $1$ [EHW24, MW19, Spi79] and [LP22] for an even stronger result for generic metrics. Numerous variants of Cartan’s result characterize other special metrics in terms of the presence of large families of special submanifolds [VLV87].
@@
Results of this kind can be found for example in Teichmüller dynamics [EFW18], in the geometry of symmetric spaces and homogeneous dynamics [BFMS21, BFMS23, BU23, LO19, MMO17, MMO22, MM22], and in algebraic geometry and the theory of variations of Hodge structure [BKU24]. We remark also that the third author obtained a result similar to Theorem 3 but under strong assumptions on the topology and curvature of $(M,g)$ [Low23] (see also [CMN22].) Studying conditions for finiteness of totally geodesic manifolds in general seems natural in the context of these prior works, see § 4.2 for a conjecture generalizing Theorem 1 with this context in mind. In Proposition 4.1.3 we construct, for each dimension $n\geq 3$ and each positive integer $k$, non-hyperbolic negatively curved analytic metrics $g_{k}$ on a fixed closed manifold $M$ that have exactly $k$ totally geodesic hypersurfaces.
The final motivation and context for our result comes from a growing literature that extends both results and techniques from homogeneous dynamics to more general smooth dynamical settings, see [Fis23, Section 1] for a brief discussion of this in the context of the study of invariant measures. A novel aspect of our proofs in this context are that they are mainly topological and not ergodic theoretic. This enables us to also prove:
@@
At the moment we can obtain a similar result in all dimensions only by assuming in addition that $M$ contains one closed totally geodesic immersed hypersurface or by assuming some curvature pinching to use the results of Cekic, Lefeuvre, Moroianu and Semmelmann [CLMS23]. A general conjecture inspired by this result is also contained in § 4.2.
@@
Our proof is based strongly on the fact that the set of hyperplanes in $TM$ tangent to totally geodesics submanifolds is a closed analytic set. This follows from the fact that the second fundamental form is an analytic function and totally geodesic submanifolds are characterized by vanishing of the second fundamental form, see § 3.2.1. We can view this closed analytic set as living in $HM=\operatorname{{Gr}}_{n-1}(TM)$ the Grassman bundle of hyperplanes in $TM$. The goal is to show that this analytic set contains an open set, so is all of $HM$, at which point we are done by Theorem 2. A main input from the theory of analytic sets is that since this set contains infinitely many totally geodesic manifolds of dimension $n-1$, it must be of dimension at least $n$. In homogeneous dynamics one typically enlarges or regularizes an invariant set or measure by proving “extra invariance” under more group directions, here we use this jump in dimension from the theory of analytic sets instead. While the argument is simple, it is also, to the best of our knowledge, novel and introduces a technique that should have other applications.
@@
The rest of the proof involves exploiting this dimension jump. This is done in two steps, the first using the dynamics of the geodesic flow, the next using the dynamics of the frame flow. At the first step we use the dynamics of the geodesic flow to show that each unit tangent vector is contained in some totally geodesic hypersurface, see § 3.2.4. This is done by showing that the set of such vectors intersect either the unstable or stable manifold in an open set, which then implies they are dense under either the forward or backward flow.
@@
In odd dimensions, the frame flow step follows from a theorem of Brin and Brin–Gromov. If $UM$ denotes the unit tangent bundle of $M$ and $PM$ denotes the bundle of orthogonal frames over $M$, the main results of [Bri75, BG80] imply that a closed invariant sets in $PM$ covering all of $UM$ is all of $PM$.
@@
In even dimensions the frame flow step is more novel and requires an analysis of closed invariant sets in $PM$ projecting to all of $UM$. For this we use that the presence of even one closed immersed totally geodesic submanifold $N$ in $M$ provides a great deal of transitivity of the frame flow on $PM$ using the Brin–Gromov result showing transitivity of the frame flow on $PN$. While this does not suffice to show transitivity on $PM$, it allows us to make a geometric argument reducing the transitivity on $HM$ to the fact that there are no non-vanishing vector fields on even dimensional spheres, see § 3.3.4.
@@
We note that A. Zeghib proved in [Zeg91, Thm. C] the analogue of Theorem 3 and Theorem 4 when the totally geodesic hypersurfaces are embedded, i.e. without self-intersections. The methods are completely different from ours and based on a lower bound of the volume of a neighborhood of an embedded hypersurface. Zeghib’s results hold in the same form for constant and variable curvature and so are very different in nature from ours which characterize constant curvature.
@@
We also note that the theory of analytic sets has been used on occasion to study fixed point sets in actions of large groups [Bon92, FS99]. However, our use of it to study non-fixed sets invariant under flows seems novel.
@@
Finally the techniques used in this paper are more or less entirely disjoint from those used in [BFMS21, BFMS23]. There is a common underlying idea of equidistribution but both how it is proven and how it is used are entirely different.
@@ 2. Preliminaries from dynamics
In § 2.1 we recall standard facts on Anosov flows. Next, in § 2.2 we take the opportunity to revisit some constructions of Brin [Bri75] regarding extensions of Anosov flows by compact groups. We take the point of view described in [EFW18, §2] and [Fil24, §5.2]. The picture is controlled by a family of closed subgroups of the compact group giving the extension. A group is naturally associated to each point on the base of the flow, and for distinct points the groups are isomorphic, but not canonically.
@@ 2.1.1. Setup
When $M$ is a compact manifold with negative sectional curvature, the geodesic flow on its unit tangent bundle $UM$ is Anosov ([KH95, §17.6]).
@@ 2.1.2. Assumptions for stable/unstable manifolds
All that we need for our arguments is the existence of stable and unstable foliations, as described in § 2.1.3, together with their standard properties as described below. See also [BV08] for a recent treatment which shows that the invariant manifolds are real-analytic when the dynamics is.
@@
We use the norms from the complexification since, in general, placing a topology or a system of seminorms on real-analytic functions is rather subtle, see [KP02, §2.6].
@@ 2.1.3. Stable/Unstable manifolds
With these assumptions, it follows from standard hyperbolic theory, as treated for example in [BP07, §7] or [Pes04, §4], that we have a foliation of $UM$ by immersed stable manifolds, with the leaf containing $x$ denoted $\mathcal{W}^{s}[x]$ and the local manifold $\mathcal{W}^{s}_{loc}[x]\subset\mathcal{W}^{s}[x]\cap B(x,r(x))$ denoting the connected component containing $x$.
@@ 2.1.4. Hopf argument
The volume measure preserved by $g_{t}$ is ergodic, as follows from the Hopf argument, see e.g. [BP07, Thm. 9.1.1]. In particular, by the Birkhoff ergodic theorem the set of points $D\subset UM$ for which both the forward and backward $g_{t}$-orbits are dense is a set of full measure.
@@
In the case of negatively curved metrics, one can also verify density of $D^{+/-}$ from topological transitivity of the geodesic flow. Transitivity of the flow can be proven geometrically, see e.g. [Ebe72].
@@ 2.1.5. Center-stable projection
A submanifold $N\subset W$ with $\dim N=\dim\mathcal{W}^{u}$ will be called transversal to the $\mathcal{W}^{cs}$-foliation if $T_{x}N\oplus W^{cs}(x)=T_{x}UM$. In this case $\pi^{cs}_{u}$ is a homeomorphism between $N$ and its image (see also [BP07, p. 236, Thm. 8.4.3]).
@@ 2.1.7. Recurrence and ergodicity in the noncompact case
Suppose $UM$ is the unit tangent bundle of a complete, finite volume Riemannian manifold $M$ which is not compact. Suppose furthermore that all the sectional curvatures of $M$ lie in $\left[-C,\tfrac{-1}{C}\right]$ for some $C>0$. A straightforward adaptation of [FLMS21, Lemma 5.13] implies that there exists a compact set $M_{th}\subset M$ such that every totally geodesic hypersurface intersects $M_{th}$. Indeed, the necessary statement about virtual nilpotency of the fundamental groups of the cusps follows from the Margulis lemma, see e.g. [BGS85, §8]. The corresponding thick-thin decomposition is constructed in [BGS85, §10.5].
For our proof, we also need to know the geodesic flow on $UM$ is Anosov and ergodic. In finite volume this follows from requiring both that all the sectional curvatures of $M$ lie in $\left[-C,\tfrac{-1}{C}\right]$ for some $C>0$ and that all derivatives of the curvature tensor are uniformly bounded. This is explained by Brin on page $82$ of his exposition of the proof in an appendix to [Bal95].
@@ 2.2.7. Stable manifolds on the principal bundle
where $T^{v}P\subset TP$ denotes the vertical subbundle (the kernel under the projection $P\to U$), and $\widetilde{W}^{c}(p)$ is the direction of the vector field generating $G_{t}$. Furthermore, the flow $G_{t}$ on $P$ also admits stable/unstable manifolds $\widetilde{\mathcal{W}}^{s/u}[p]$ with the same defining properties as in § 2.1.3: $p^{\prime}\in\widetilde{\mathcal{W}}^{s}[p]$ if and only if $\operatorname{{dist}}(G_{t}p^{\prime},G_{t}p)\xrightarrow{t\to+\infty}0$ and $T_{p}\widetilde{\mathcal{W}}^{s}[p]=\widetilde{W}^{s}(p)$. If $p\in P$ projects to $u\in U$, then $\widetilde{\mathcal{W}}^{s}_{loc}[p]$ projects bijectively to $\mathcal{W}^{s}_{loc}[u]$. For the existence and properties of stable/unstable foliations in this context, we refer to [Pes04, §4].
@@ 2.2.14. Minimal sets
The relatively minimal set $X$ gives a map $[X]\colon U\to P/H$, which is continuous because $X$ is closed and the fibers of the principal bundle are compact. This map, or the corresponding $H$-principal subset $X\subset P$, is the desired reduction. ∎
@@
$cl\colon\left.\raisebox{1.00006pt}{$P$}\middle/\raisebox{-1.00006pt}{$K_{1}$}\right.\to\raisebox{-1.00006pt}{$H$}\backslash\raisebox{1.00006pt}{$K$}/\raisebox{-1.00006pt}{$K_{1}$}$
@@
    After passing to a finite cover of the base manifold, we can ensure that the transitivity group is connected, see [CLMS23, Lemma 3.3]. We do so in the applications in Section 3.

### v3

@@ 1. Introduction
In the above statement, “hyperbolic” refers to a Riemannian manifold with constant strictly negative sectional curvature. Equivalently, up to rescaling the Riemannian metric, $M$ is isometric to $\mathbb{H}^{n}/\Gamma$ where $\mathbb{H}^{n}$ is hyperbolic $n$-space, and $\Gamma$ is a lattice which is moreover arithmetic (see [Margulis_Discrete-subgroups-of-semisimple-Lie-groups1991, Ch. IX] for the notion of arithmetic lattice).
@@
    Our proof works under the more general condition that the geodesic flow of $(M,g)$ is Anosov, see e.g. [Eberlein1973_When-is-a-geodesic-flow-of-Anosov-type-III, Thm. 3.2] for equivalent characterizations of this property. Our proof also applies to $M$ of finite volume instead of compact, with some additional technical assumptions spelled out in § 2.1.2 and § 2.1.7.
@@
    The arithmeticity of $M$ is a consequence of the prior conclusions and [BFMS1, Thm. 1.1]. See also [MohammadiMargulis2022_Arithmeticity-of-hyperbolic-3-manifolds-containing-infinitely-many-totally, Thm. 1.1] for the case $n=3$.
@@
    Arithmetic hyperbolic manifolds of the first type have infinitely many closed totally geodesic immersed hypersurfaces. All other arithmetic hyperbolic manifolds do not. See [belolipetsky2021subspace, Corollary 5.12] and the remark following [belolipetsky2021subspace, Theorem 1.2].
  4. (iv)
    It would be interesting to weaken the assumption of analyticity to smoothness, or relax the Anosov assumption to the manifold being of rank $1$ and nonpositively curved (see [knieper98] for this last condition). In work in preparation, we will address the rank $1$ case, where without analyticity of the metric the result does not hold. See also [BallmannGromovSchroeder1985_Manifolds-of-nonpositive-curvature, Lecture III] for other results that hold in the analytic but not smooth setting of nonpositively curved manifolds.
    Additionally, it is worth noting that all other currently available finiteness results that motivate this work are in the homogeneous, or complex-analytic and algebraic settings, which are substantially more constrained than our case. For example, they only apply to geometric settings that only depend on finite-dimensional parameters.
@@
See [Cartan] and [Dajczer1990_Submanifolds-and-isometric-immersions, Thm. 1.8] for modern presentations of the proof. The axiom of planes, which requires that through any three points in a Riemannian manifold passes a totally geodesic surface, was originally introduced by Riemann [riemann1854uber].
Theorem 2 is not only a significant motivation for our work but also a key step in our proof. We note that in this context there is no version of Theorem 1 for manifolds of positive curvature or for general non-compact manifolds, see § 4.1. We also remark that there is always an open dense set of metrics on a manifold for which there are no totally geodesic submanifolds of dimension greater than $1$ [ElHWilhelm, MurphyWilhelm, Spivak] and [LytchakPetrunin] for an even stronger result for generic metrics. Numerous variants of Cartan’s result characterize other special metrics in terms of the presence of large families of special submanifolds [survey_tot_geod_submanifolds].
@@
Results of this kind can be found for example in Teichmüller dynamics [EskinFilipWright2018_The-algebraic-hull-of-the-Kontsevich-Zorich-cocycle], in the geometry of symmetric spaces and homogeneous dynamics [BFMS1, BFMS2, BaldiUllmo, LeeOh, MMO1, MMO2, MohammadiMargulis2022_Arithmeticity-of-hyperbolic-3-manifolds-containing-infinitely-many-totally], and in algebraic geometry and the theory of variations of Hodge structure [BaldiKlinglerUllmo]. We remark also that the third author obtained a result similar to Theorem 3 but under strong assumptions on the topology and curvature of $(M,g)$ [Lowe] (see also [cmn22].) Studying conditions for finiteness of totally geodesic manifolds in general seems natural in the context of these prior works, see § 4.2 for a conjecture generalizing Theorem 1 with this context in mind. In Proposition 4.1.3 we construct, for each dimension $n\geq 3$ and each positive integer $k$, non-hyperbolic negatively curved analytic metrics $g_{k}$ on a fixed closed manifold $M$ that have exactly $k$ totally geodesic hypersurfaces.
The final motivation and context for our result comes from a growing literature that extends both results and techniques from homogeneous dynamics to more general smooth dynamical settings, see [Fisher-ICM, Section 1] for a brief discussion of this in the context of the study of invariant measures. A novel aspect of our proofs in this context are that they are mainly topological and not ergodic theoretic. This enables us to also prove:
@@
At the moment we can obtain a similar result in all dimensions only by assuming in addition that $M$ contains one closed totally geodesic immersed hypersurface or by assuming some curvature pinching to use the results of Cekic, Lefeuvre, Moroianu and Semmelmann [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds]. A general conjecture inspired by this result is also contained in § 4.2.
@@
Our proof is based strongly on the fact that the set of hyperplanes in $TM$ tangent to totally geodesics submanifolds is a closed analytic set. This follows from the fact that the second fundamental form is an analytic function and totally geodesic submanifolds are characterized by vanishing of the second fundamental form, see § 3.2.1. We can view this closed analytic set as living in $HM=\operatorname{{Gr}}_{n-1}(TM)$ the Grassman bundle of hyperplanes in $TM$. The goal is to show that this analytic set contains an open set, so is all of $HM$, at which point we are done by Theorem 2. A main input from the theory of analytic sets is that since this set contains infinitely many totally geodesic manifolds of dimension n1n-1, it must be of dimension at least $n$. In homogeneous dynamics one typically enlarges or regularizes an invariant set or measure by proving “extra invariance” under more group directions, here we use this jump in dimension from the theory of analytic sets instead. While the argument is simple, it is also, to the best of our knowledge, novel and introduces a technique that should have other applications.
@@
The rest of the proof involves exploiting this dimension jump. This is done in two steps, the first using the dynamics of the geodesic flow, the next using the dynamics of the frame flow. At the first step we use the dynamics of the geodesic flow to show that each unit tangent vector is contained in some totally geodesic hypersurface, see § 3.2.4. This is done by showing that the set of such vectors that intersect either the unstable or stable manifold in an open set, which then implies they are dense under either the forward or backward flow.
@@
In odd dimensions, the frame flow step follows from a theorem of Brin and Brin–Gromov. If $UM$ denotes the unit tangent bundle of $M$ and $PM$ denotes the bundle of orthogonal frames over $M$, the main results of [Brin1975_The-topology-of-group-extensions-of-C-systems, BrinGromov1980_On-the-ergodicity-of-frame-flows] imply that a closed invariant set in $PM$ covering all of $UM$ is all of $PM$.
@@
In even dimensions the frame flow step is more novel and requires an analysis of closed invariant sets in $PM$ projecting to all of $UM$. For this we use that the presence of even one closed immersed totally geodesic hypersurface $N$ in $M$ provides a great deal of transitivity of the frame flow on $PM$ using the Brin–Gromov result showing transitivity of the frame flow on $PN$. While this does not suffice to show transitivity on $PM$, it allows us to make a geometric argument reducing the transitivity on $HM$ to the fact that there are no non-vanishing vector fields on even dimensional spheres, see § 3.3.4.
@@
We note that A. Zeghib proved in [Zeghib_Laminations-et-hypersurfaces-geodesiques-des-varietes1991, Thm. C] the analogue of Theorem 3 and Theorem 4 when the totally geodesic hypersurfaces are embedded, i.e. without self-intersections. The methods are completely different from ours and based on a lower bound of the volume of a neighborhood of an embedded hypersurface. Zeghib’s results hold in the same form for constant and variable curvature and so are very different in nature from ours which characterize constant curvature.
@@
We also note that the theory of analytic sets has been used on occasion to study fixed point sets in actions of large groups [Bonatti, FarbShalen]. However, our use of it to study non-fixed sets invariant under flows seems novel.
@@
Finally the techniques used in this paper are more or less entirely disjoint from those used in [BFMS1, BFMS2]. There is a common underlying idea of equidistribution but both how it is proven and how it is used are entirely different.
@@ 2. Preliminaries from dynamics
In § 2.1 we recall standard facts on Anosov flows. Next, in § 2.2 we take the opportunity to revisit some constructions of Brin [Brin1975_The-topology-of-group-extensions-of-C-systems] regarding extensions of Anosov flows by compact groups. We take the point of view described in [EskinFilipWright2018_The-algebraic-hull-of-the-Kontsevich-Zorich-cocycle, §2] and [Filip_Translation-surfaces:-Dynamics-and-Hodge-theory, §5.2]. The picture is controlled by a family of closed subgroups of the compact group giving the extension. A group is naturally associated to each point on the base of the flow, and for distinct points the groups are isomorphic, but not canonically.
@@ 2.1.1. Setup
When $M$ is a compact manifold with negative sectional curvature, the geodesic flow on its unit tangent bundle $UM$ is Anosov ([KatokHasselblatt1995_Introduction-to-the-modern-theory-of-dynamical-systems, §17.6]).
@@ 2.1.2. Assumptions for stable/unstable manifolds
All that we need for our arguments is the existence of stable and unstable foliations, as described in § 2.1.3, together with their standard properties as described below. See also [BarreiraValls2008_Analytic-invariant-manifolds-for-sequences-of-diffeomorphisms] for a recent treatment which shows that the invariant manifolds are real-analytic when the dynamics is.
@@
We use the norms from the complexification since, in general, placing a topology or a system of seminorms on real-analytic functions is rather subtle, see [krantz2002primer, §2.6].
@@ 2.1.3. Stable/Unstable manifolds
With these assumptions, it follows from standard hyperbolic theory, as treated for example in [BarreiraPesin2007_Nonuniform-hyperbolicity, §7] or [Pesin2004_Lectures-on-partial-hyperbolicity-and-stable-ergodicity, §4], that we have a foliation of $UM$ by immersed stable manifolds, with the leaf containing $x$ denoted $\mathcal{W}^{s}[x]$ and the local manifold $\mathcal{W}^{s}_{loc}[x]\subset\mathcal{W}^{s}[x]\cap B(x,r(x))$ denoting the connected component containing $x$.
@@ 2.1.4. Hopf argument
The volume measure preserved by $g_{t}$ is ergodic, as follows from the Hopf argument, see e.g. [BarreiraPesin2007_Nonuniform-hyperbolicity, Thm. 9.1.1]. In particular, by the Birkhoff ergodic theorem the set of points $D\subset UM$ for which both the forward and backward $g_{t}$-orbits are dense is a set of full measure.
@@
In the case of negatively curved metrics, one can also verify density of $D^{+/-}$ from topological transitivity of the geodesic flow. Transitivity of the flow can be proven geometrically, see e.g. [Eberlein].
@@ 2.1.5. Center-stable projection
A submanifold $N\subset W$ with $\dim N=\dim\mathcal{W}^{u}$ will be called transversal to the $\mathcal{W}^{cs}$-foliation if $T_{x}N\oplus W^{cs}(x)=T_{x}UM$. In this case $\pi^{cs}_{u}$ is a homeomorphism between $N$ and its image (see also [BarreiraPesin2007_Nonuniform-hyperbolicity, p. 236, Thm. 8.4.3]).
@@ 2.1.7. Recurrence and ergodicity in the noncompact case
Suppose $UM$ is the unit tangent bundle of a complete, finite volume Riemannian manifold $M$ which is not compact. Suppose furthermore that all the sectional curvatures of $M$ lie in $\left[-C,\tfrac{-1}{C}\right]$ for some $C>0$. A straightforward adaptation of [FisherLafont_Finiteness-of-maximal-geodesic-submanifolds2021, Lemma 5.13] implies that there exists a compact set $M_{th}\subset M$ such that every totally geodesic hypersurface intersects $M_{th}$. Indeed, the necessary statement about virtual nilpotency of the fundamental groups of the cusps follows from the Margulis lemma, see e.g. [BallmannGromovSchroeder1985_Manifolds-of-nonpositive-curvature, §8]. The corresponding thick-thin decomposition is constructed in [BallmannGromovSchroeder1985_Manifolds-of-nonpositive-curvature, §10.5].
For our proof, we also need to know the geodesic flow on $UM$ is Anosov and ergodic. In finite volume this follows from requiring both that all the sectional curvatures of $M$ lie in $\left[-C,\tfrac{-1}{C}\right]$ for some $C>0$ and that all derivatives of the curvature tensor are uniformly bounded. This is explained by Brin on page $82$ of his exposition of the proof in an appendix to [BallmannLectures].
@@ 2.2.7. Stable manifolds on the principal bundle
where $T^{v}P\subset TP$ denotes the vertical subbundle (the kernel under the projection $P\to U$), and $\widetilde{W}^{c}(p)$ is the direction of the vector field generating $G_{t}$. Furthermore, the flow $G_{t}$ on $P$ also admits stable/unstable manifolds $\widetilde{\mathcal{W}}^{s/u}[p]$ with the same defining properties as in § 2.1.3: $p^{\prime}\in\widetilde{\mathcal{W}}^{s}[p]$ if and only if $\operatorname{{dist}}(G_{t}p^{\prime},G_{t}p)\xrightarrow{t\to+\infty}0$ and $T_{p}\widetilde{\mathcal{W}}^{s}[p]=\widetilde{W}^{s}(p)$. If $p\in P$ projects to $u\in U$, then $\widetilde{\mathcal{W}}^{s}_{loc}[p]$ projects bijectively to $\mathcal{W}^{s}_{loc}[u]$. For the existence and properties of stable/unstable foliations in this context, we refer to [Pesin2004_Lectures-on-partial-hyperbolicity-and-stable-ergodicity, §4].
@@ 2.2.14. Minimal sets
The relatively minimal set $X$ gives a map $X]\colon U\to P/H$, which is continuous because $X$ is closed and the fibers of the principal bundle are compact. This map, or the corresponding $H$-principal subset $X\subset P$, is the desired reduction. ∎
@@
[…]
@@
    After passing to a finite cover of the base manifold, we can ensure that the transitivity group is connected, see [CekicLefeuvreMoroianu2023_On-the-ergodicity-of-the-frame-flow-on-even-dimensional-manifolds, Lemma 3.3]. We do so in the applications in Section 3.